# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ 7.1 Wildfire risk rating and probability calibration
Although Jain et al. (2020); Zacharakis and Tsihrintzis (2023b) systematically reviewed the application of various machine learning methods in wildfire risk research, they did not provide a description of the post-processing techniques required to convert the predictions of machine learning models into actual wildfire risk. Directly using the output of these models as the probability of wildfire risk may lack theoretical foundation and stability. This is because wildfire historical data sets typically suffer from severe class imbalance, where the spatial and temporal scope of fire occurrences is considerably smaller compared to non-fire events. To expedite model convergence, many machine learning datasets are often constructed with a balanced class ratio, typically 1:1, 1:1.5, or 1:2 (Huot et al., 2021; Kondylatos et al., 2022) between fire and non-fire instances. This may lead to an overestimation of false positives during model evaluation (Bakke et al., 2023). However, actual wildfire risk probabilities are usually much lower. For example, in wildfire risk assessments for the Mediterranean region and the United States, Ager et al. (2014a) and Preisler et al. (2004) used probabilities in the thousandths range, respectively.
Using the probability output of machine learning models as wildfire risk could be misleading to non-experts. Therefore, many studies convert continuous wildfire risk probabilities into risk levels to support wildfire management, such as conveying public warnings and guiding fire suppression measures. Among the methods for risk level classification, the Jenks natural breaks classification (Jenks and Caspall, 1971) is most commonly used. For instance, Xie et al. (2022) trained an ensemble machine learning model using binary wildfire ignition records and then classified the predicted probabilities into five groups using the Jenks method. Similarly, Iban and Sekertekin (2022) and Moayedi and Khasmakhi (2023) also classified the hazard maps generated by machine learning models into five categories using the Jenks method. Along the same lines, Chen et al. (2024c) used the Jenks method to divide the estimated potential rate of spread and fire radiative power into five levels for wildfire risk assessment.
@@
Furthermore, some methods aim to recalibrate probability outputs to obtain actual wildfire risk probabilities. For instance, Pelletier et al. (2023) constructed a semi-balanced dataset for model training, in which the numbers of burned and never-burned sites were 3,268 and 60,146, respectively. They then used the method proposed by Elkan (2001) (Eq. (9)) to convert the wildfire probabilities predicted by a time series XGBoost model into actual probability values:
@@
where $p^{\prime}$ represents the revised probability for a prediction, $BR1$ stands for the base rate of the predicted probability estimates (the average of predicted probabilities for that class), and $BR2$ denotes the base rate of the actual probability estimates derived from the target population. Here, the target population refers to the average probability of fires across all observations within the training dataset, including the unburned observations that were not utilized in model training. Similarly, Phelps and Woolford (2021) used the function proposed by Dal Pozzolo et al. (2015) to correct the neural network’s probability output:
@@
where $p_{k}$ and $\gamma_{k}$ represent the wildfire probabilities modeled from the original and sampled distributions, respectively, and $\pi$ denotes the proportion of non-fire observations sampled.
@@ 7.2.1 Recurrent Neural Networks
As discussed in previous sections, the occurrence and spread of wildfires are influenced by various factors, including meteorological conditions, fuel status, and historical wildfire data. These factors exhibit cumulative effects. Therefore, utilizing time series data enables models to capture the gradual changes in influencing factors, thus facilitating the prediction of future wildfire risks. The most classical methods in time series prediction tasks are RNNs, which include RNN, LSTM, and GRU.
RNN is a basic neural network structure with recurrent connections, designed to process sequential data. As shown in Fig. 6, it integrates input from the current time step with output from the previous time step to model and predict sequence data. A simple RNN structure consists of an input layer, hidden layer, and output layer. Input data are time series data, i.e., feature vectors with dimensions $f\times d$. The hidden layer includes several hidden units, where each hidden unit outputs $\mathbb{h}_{t-m}$, the ’memory’ at any intermediate time $t-m(m<n)$, as a weighted sum of its previous ’memory’ and current time step’s input features:
Figure 6: The architecture of RNN where (a) and (b) represent the overall architecture and detailed hiddenlayer calculation.
@@
Classic RNN networks and their variants have been employed in wildfire risk prediction. For instance, Cheng and Wang (2008) designed a spatiotemporal Elman RNN-based prediction framework integrating historical wildfire and weather data to predict the annual average burned area in Canadian forests. Experiments demonstrated a prediction error of less than 0.5ha on areas of 500ha, validating the feasibility of using RNN for fire prediction. Similarly, a comparative study of ten artificial neural networks (ANNs) based on meteorological data for predicting wildfire incidence in Australia found that Elman RNN exhibited the best performance, with average accuracy, sensitivity, and specificity exceeding 93% (Dutta et al., 2013). Furthermore, a network composed of a dynamic autoencoder and RNN was used to predict the next month’s burned area in five U.S. regions, where the dynamic autoencoder was employed to transform multidimensional time series data into RNN input features, and dual decoders were used for predicting fire spot likelihood over different time spans, i.e., one week and one month. Comparative experiments showed that their method was slightly superior to comparative methods, generative network, and GRU (Chavalithumrong et al., 2021).
@@
While Jain et al. (2020) and others have highlighted the limited use of classical RNNs, our findings are consistent with these observations. The underutilization of RNNs may be attributed to their difficulties in handling long sequence data, where they are prone to gradient vanishing or exploding, making it challenging to capture long-term dependencies (Hochreiter, 1997; Gers et al., 2000). To address these issues, LSTM networks were developed. As illustrated in Fig. 7, LSTMs incorporate three gating units: input gate, forget gate, and output gate, which control the flow and retention of information. These gates have trainable weights and decide whether to pass, forget, or output information based on the current input and the output from the previous time steps. This gating mechanism effectively resolves the issues of vanishing and exploding gradients, thereby enabling better handling of long sequences and capturing of long-term dependencies. Consequently, LSTMs are more prevalently applied in wildfire risk prediction than classical RNNs. The specific configuration of an LSTM cell is depicted in Fig. 7b and Eq.13 (Monner and Reggia, 2012; Liang et al., 2019).
Figure 7: Schematic diagram of the LSTM architecture illustrating the flow of data through input, forget, and output gates.
  • 1.
    For the forget gate ($f_{t}$), the output from the hidden layer of the previous timestep $h_{t-(m+1)}$ is concatenated with the input features at the current timestep $x_{t-m}$, then multiplied by the forget gate weight matrix $W_{f}$ and added to the forget gate bias $b_{f}$. This sum is then passed through a sigmoid activation function $\sigma$, resulting in the output of the forget gate.
  • 2.
    The input gate has two branches. The first branch ($i_{t}$) is similar to the forget gate, involving concatenation followed by multiplication with the input gate weight matrix $W_{i}$ and addition of the bias $b_{i}$. The second branch uses a weight matrix $W_{c}$ that represents candidate cell states to multiply with the concatenated result, added to the bias $b_{c}$, and activated using the hyperbolic tangent function $\tanh$. The outputs of these two branches are multiplied together and added to the product of the forget gate output and the output of the input gate from the previous timestep ($c_{t-(m+1)}$) to produce the output for the current timestep ($c_{t-m}$).
  • 3.
    Similarly, the output gate includes two branches. The first branch multiplies the concatenated input by the output gate weight matrix $W_{o}$, adds the bias $b_{o}$, and activates it. For the second branch, the output of the input gate at the current timestep ($c_{t}$) is activated using $\tanh$. The final output of the hidden layer at the current timestep is obtained by multiplying the results of these two branches.
$f_{t-m}=$ $\sigma\left(W_{f}\cdot\left[h_{t-(m+1)},x_{t-m}\right]+b_{f}\right)$ (13)
$i_{t-m}=$ $\sigma\left(W_{i}\cdot\left[h_{t-(m+1)},x_{t-m}\right]+b_{i}\right)$ (14)
$c_{t-m}=$ $f_{t}\odot c_{t-1}+i_{t}\odot\tanh\left(W_{c}\cdot\left[h_{t-1},x_{t}\right]+b_{c}\right)$ (15)
$o_{t-m}=$ $\sigma\left(W_{o}\cdot\left[h_{t-(m+1)},x_{t-m}\right]+b_{o}\right)$ (16)
$h_{t-m}=$ $o_{t-m}\odot\tanh(c_{t-m})$ (17)
@@
Moreover, some studies use only historical wildfire occurrence data combined with LSTM to predict future regional wildfire risks. For example, Kadir et al. (2023) used historical MODIS wildfire data combined with LSTM to predict the spatiotemporal distribution of wildfires in Indonesia for the following year, with a success rate exceeding 90% and an error rate of only 6.94%. Notably, to reduce the amount of data, this study grouped several days’ worth of data into single days for training and testing. Furthermore, Hu et al. (2023) combined LSTM with autoencoders, using historical wildfire data from 1992-2018 to predict the level of wildfire occurrence in high-risk areas such as California. The autoencoder was employed to generalize wildfire events from the overall dataset, enhancing the accuracy of anomaly event predictions through reconstruction errors.
Similar to the LSTM, the GRU addresses the challenges of gradient vanishing and explosion inherent in traditional RNNs and involving the gate mechanisms. GRU reduces the number of gating units found in LSTM, containing only a reset gate and an update gate, as shown in Eq.18, which simplifies the model architecture and enhances computational efficiency (Chung et al., 2014). The structure of the GRU model is depicted in Fig. 8. Despite its simpler configuration, GRU retains the capability to model long-term dependencies with performance comparable to that of LSTM (Jin et al., 2020a).
$r_{t}=\sigma(W_{r}[h_{t-1},x_{t}]+b_{r})$ (18)
$z_{t}=\sigma(W_{z}[h_{t-1},x_{t}]+b_{z})$ (19)
$\tilde{h}_{t}=\tanh(W_{h}[r_{t}h_{t-1},x_{t}]+b_{h})$ (20)
$h_{t}=(1-z_{t})h_{t-1}+z_{t}\tilde{h}_{t}$ (21)
Figure 8: Schematic representation of the GRU architecture highlighting its reset and update gates.
Studies such as Dzulhijjah et al. (2023) have compared the performance of LSTM, BiLSTM, CNN-LSTM, Stacked-LSTM, and GRU in predicting wildfire risk in the Indonesian region of Kalimantan, finding that GRU slightly outperforms the other models, though the advantage is not significant. In contrast, Gopu et al. (2023) observed better results with LSTM in a comparative study of wildfire prediction in the Montesinho Natural Park, Portugal. Additionally, Chavalithumrong et al. (2021) found that models integrating a dynamic autoencoder with RNN outperformed standard GRU models, suggesting that enhanced RNN architectures can offer superior performance.
@@ 7.2.2 Transformers for Wildfire Risk Prediction
where $d_{k}$ represents the dimensionality of the key. The architecture of a Transformer, as illustrated in Fig. 9, comprises encoders, decoders, positional encoding, and multiple attention mechanisms.
@@
Transformers have been adapted for several predictive tasks in wildfire risk prediction. For example, Miao et al. (2023) developed a Transformer model with a window-based attention mechanism to predict forest fire risks in Chongli District, Beijing, China, using time-series data of meteorological, topographical, vegetation, and anthropogenic factors. The window mechanism confines attention to a fixed-size window centered on each element, reducing computational complexity and memory usage. Similarly, Cao et al. (2024) employed a Transformer to predict next-day wildfire risks in Quanzhou County, Guangxi Province, China, based on three days of meteorological, topographical, and human activity data. The model’s performance, analyzed using IGR, outperformed LSTM, RNN, and SVM in terms of generalization, noise resistance, and overfitting mitigation.
Figure 9: Schematic diagram of the Transformer architecture, highlighting its main components.
@@ 7.3.1 Convolutional Neural Networks
CNNs are foundational network structures in deep learning, designed to extract local spatial features through convolution operations, making them well-suited for capturing the spatial information necessary for wildfire risk prediction, as shown in Fig.10. For example, Zhang et al. (2019) employed a modified AlexNet (Krizhevsky et al., 2012) using annual fire point maps and springtime mean influencing factors for Yunnan Province, China, to classify image patches as wildfire-prone areas. The model outputs a probability map of wildfire risk through softmax mapping in the final CNN layer. When comparing the CNN model’s performance with RF, SVM, MLP, and kernel logistic regression models, the authors found that CNN achieved the highest overall accuracy. They also analyzed the contributions of various wildfire drivers using the information gain ratio, identifying temperature, wind speed, surface roughness, precipitation, and elevation as the most influential factors.
@@
Building upon this, the authors extended the modified AlexNet to generate a global quarterly wildfire susceptibility map (Zhang et al., 2021). They compared its performance with MLP-2D, CNN-1D, and MLP-1D models, finding that CNN-2D and MLP-1D models outperformed CNN-1D and MLP-2D. Additionally, they applied an improved permutation importance (PI) method and partial dependence plots (PDP) to investigate the interpretability of the CNN-2D model. The PI analysis highlighted that maximum monthly temperature, soil temperature, NDVI, and soil moisture were significant factors affecting the model. Furthermore, PDP analysis revealed a negative relationship between cumulative precipitation and wildfire probability, while maximum monthly temperature and soil temperature positively influenced wildfire occurrence.
@@
Moreover, using a similar CNN approach, they predicted the spatiotemporal variation of wildfire risk under different climate scenarios, discovering that global warming would lead to an increase in burned areas and a northward shift of wildfire-prone regions (Zhang et al., 2024). In another study, Kanwal et al. (2023) explored CNN-1D and CNN-2D classification models along with other machine learning algorithms for evaluating seasonal wildfire risks in Pakistan, finding that CNN-2D outperformed CNN-1D and other machine learning algorithms.
@@
In addition, some studies have employed deeper or structurally optimized CNN networks. For instance, Oak et al. (2024) utilized VGG16 with ReLU, BatchNorm, Dropout, and a sigmoid layer to predict wildfire susceptibility in Quebec, Canada. Comparative experiments demonstrated that this model outperformed other architectures like Xception, VGG16, ResNet50, and Inception V3. Similarly, Nur et al. (2022) applied a meta-heuristic optimization algorithm to fine-tune the hyperparameters of a CNN model for wildfire susceptibility mapping in the Plumas National Forest, USA. Their IGR analysis revealed that land use, drought index, and maximum temperature had the greatest impact on wildfire susceptibility, and the model’s performance improved significantly after optimization. Likewise, (You et al., 2023) used particle swarm optimization to refine the structure and parameters of a CNN classification model for assessing annual forest fire risk across China, showing that this approach outperformed both standard CNN and traditional machine learning methods. Using SHAP for interpretability analysis, they identified NDVI, land cover, and temperature as the most significant factors.
Additionally, Eddin et al. (2023) introduced location information to the CNN by adding a separate convolution branch to differentiate between dynamic variables (those changing over time) and static variables (those that remain constant over the study period). To address the causal effects of static variables on dynamic ones, they incorporated a LOAN layer to adjust the activation maps in the dynamic branch based on static features. They also used sinusoidal encoding of Julian days to provide the model with explicit time information. Their experiments on the FireCube dataset, which covers parts of the Mediterranean (Prapas et al., 2022; Kondylatos et al., 2023), demonstrated that the proposed model’s predictions of $t+1$ wildfire spread outperformed baseline models, including LSTM, ConvLSTM, RF, and XGBoost. In a similar vein, Dong et al. (2024) proposed an architecture that models dynamic and static variables separately in the encoding phase, improving the accuracy in predicting fire radiative power.
@@
Beyond classification-based methods, some studies have implemented end-to-end models for higher-resolution predictions. For example, Hodges and Lattimer (2019) proposed a deep convolutional inverse graphic network comprising convolution and deconvolution layers to simulate wildfire spread over homogeneous and heterogeneous landscapes for 6-hour and 24-hour intervals, demonstrating the high efficiency of CNNs in wildfire spread prediction. Similarly, Santopaolo et al. (2021) used an end-to-end CNN image segmentation model to predict wildfire risk in Sicily, Italy, and California, USA, based on driving factors from 8 days prior.
@@
Furthermore, (Allaire et al., 2021) developed a hybrid architecture combining convolutional and fully connected layers to emulate the output of a numerical fire spread simulator based on the Rothermel model of fire propagation. Similarly, Bergado et al. (2021) designed an AllConvNet, a fully convolutional network, to predict wildfire burn probability for the next 7 days in Victoria, Australia. This model achieved higher prediction accuracy compared to logistic regression, SegNet, and MLP models. Feature extraction and normalization of logistic regression coefficients revealed that total rainfall, lightning density, and surface temperature were among the most influential variables.
@@
Additionally, Li et al. (2023c) introduced spatial and channel attention mechanisms to enhance an end-to-end CNN model’s focus on critical features while suppressing irrelevant ones for predicting the next moment’s wildfire spread. This model also incorporated the burning map from the previous time step, outperforming other CNN models and ConvLSTM. Similarly, Shadrin et al. (2024b) used the MA-Net segmentation model, which includes position attention and multi-scale attention blocks, to predict wildfire extent for the next 1-5 days. Unlike previous studies, they included short-term weather forecasts along with various wildfire drivers. Experimental results showed that their model accurately predicted wildfire spread within 3 days, with the wind components, land cover, NDVI, EVI, and LAI being the most influential factors.

### v5

@@ 6.1 Wildfire Risk Rating and Probability Calibration
Although Jain et al. (2020); Zacharakis and Tsihrintzis (2023) systematically reviewed the application of various machine learning methods in wildfire risk research, they did not provide a description of the post-processing techniques required to convert the predictions of machine learning models into actual wildfire risk. Directly using the output of these models as the probability of wildfire risk may lack theoretical foundation and stability. This is because wildfire historical data sets typically suffer from severe class imbalance, where the spatial and temporal scope of fire occurrences is considerably smaller compared to non-fire events. To expedite model convergence, many machine learning datasets are often constructed with a balanced class ratio, typically 1:1, 1:1.5, or 1:2 (Huot et al., 2021; Kondylatos et al., 2022) between fire and non-fire instances. This may lead to an overestimation of false positives during model evaluation (Phelps and Woolford, 2021b, a; Bakke et al., 2023). However, actual wildfire risk probabilities are usually much lower. For example, in wildfire risk assessments for the Mediterranean region and the United States, Ager et al. (2014a) and Preisler et al. (2004) used probabilities in the thousandths range, respectively.
Using the probability output of machine learning models as an indicator of wildfire risk could be misleading to non-experts. Therefore, many studies convert continuous wildfire risk probabilities into risk levels to support wildfire management, such as conveying public warnings and guiding fire suppression measures. Among the methods for risk level classification, the Jenks natural breaks classification (Jenks and Caspall, 1971) is most commonly used. For instance, Xie et al. (2022b) trained an ensemble machine learning model using binary wildfire ignition records and then classified the predicted probabilities into five groups using the Jenks method. Similarly, Iban and Sekertekin (2022) and Moayedi and Khasmakhi (2023) also classified the hazard maps generated by machine learning models into five categories using the Jenks method. Along the same lines, Chen et al. (2024b) used the Jenks method to divide the estimated potential rate of spread and fire radiative power into five levels for wildfire risk assessment.
@@
In comparison to methods that predict risk classifications or directly utilize confidence levels from deterministic models as risk indicators, there remains a paucity of research focused on recalibrating probability outputs to obtain actual wildfire risk probabilities through post-processing and pre-processing of model predictions and inputs. For instance, Pelletier et al. (2023) constructed a semi-balanced dataset for model training, in which the numbers of burned and never-burned sites were 3,268 and 60,146, respectively. They then used the method proposed by Elkan (2001) (Eq. (10)) to convert the wildfire probabilities predicted by a time series XGBoost model into actual probability values:
@@
where $p^{\prime}$ represents the revised probability for a prediction, $BR1$ is the base rate of the predicted probability estimates (the average of predicted probabilities for that class), and $BR2$ denotes the base rate of the actual probability estimates derived from the target population. Here, the target population refers to the average probability of fires across all observations within the training dataset, including the unburned observations that were not utilized in model training. After comparing the fire probabilities before and after calibration, they found that the calibrated probabilities more accurately reflected the true likelihood of wildfire occurrence.
Similarly, Phelps and Woolford (2021a) used the function proposed by Dal Pozzolo et al. (2015) to correct the neural network’s probability output:
@@
where $p_{k}$ and $\gamma_{k}$ represent the wildfire probabilities modeled from the original and sampled distributions, respectively, and $\pi$ denotes the proportion of non-fire observations sampled. Furthermore, this study also employed Platt Scaling and Offset methods to calibrate tree-based methods (bagged classification trees/random forest) and statistical models (logistic regression and logistic generalised additive models), respectively. The research ultimately found that uncalibrated machine learning models severely overestimate fire probabilities (e.g., predicted values several orders of magnitude higher than actual probabilities).
Caron et al. (2024) propose a preprocessing method for calibrating wildfire ignition prediction probabilities. The core concept involves transforming binary fire events (occurrence or non-occurrence) into continuous fire risk intensity curves during the preprocessing phase, then mapping these through logistic regression models to produce calibrated probability predictions. This approach addresses class imbalance issues in wildfire risk prediction while avoiding the overestimation problems caused by sampling and uncalibrated models. Specifically, the research first applies a Laplacian convolution kernel to diffuse daily fire event sequences, thereby reducing the significance of isolated fire events while enhancing risk values for dates adjacent to continuous fire events. Subsequently, the authors calculate fire influence values for a given day using multi-scale sliding windows (Eq. (12)), incorporating both short-term and long-term trends.
$\text{Influence}(J)=\sum_{t=1}^{T}\left(\frac{1}{2t+1}\sum_{i=-t}^{t}\text{Fires}_{J+i}\right)$ (12)
where, $T$ and $J$ represent the maximum time window length and the current day, respectively, while $\text{Fires}_{J+i}$ indicates the number of fire events on day $J+i$. Subsequently, $\text{Influence}(J)$ serves as input to a logistic regression model, which learns the relationship between this value and actual fire occurrences (0/1), yielding calibrated probability outputs. Evaluation of this method demonstrates excellent calibration performance. Assessment metrics include Log Loss and Expected Calibration Error (ECE). Compared to uncalibrated traditional classification models, this approach significantly reduces ECE (from approximately 10% to 2%–4%) while maintaining low Log Loss, proving that its predictions achieve a favorable balance between accuracy and reliability (Caron et al., 2024).
In addition to the aforementioned methods, Zhang et al. (2025) introduced a two-stage parametric calibration framework specifically designed to address distributional shifts in deep learning-based drought detection, which is conceptually transferable to wildfire probability calibration. This two-step process corrects overconfident predictions by explicitly addressing spatial heterogeneity and sampling-induced class imbalance. In the first stage, they divided the study area into $N$ climate zones and fitted zone-specific logistic calibration functions to mitigate label prior shift. For each zone $K_{i}$, the calibrated probability $\hat{p}$ is computed as:
$\hat{p}=\frac{1}{1+\exp(-a_{i}\cdot\log\frac{p}{1-p}-b_{i})},$ (13)
where $p$ is the original model output (logit), and $(a_{i},b_{i})$ are the learned parameters optimized by minimizing negative log-likelihood on zone-specific training samples. In the second stage, to correct likelihood shift across unseen validation data, they introduced a bias-corrected transformation:
$\tilde{p}=\frac{\beta\hat{p}}{\beta\hat{p}-\hat{p}+1},$ (14)
where $\beta$ is tuned by minimizing the ECE over binned confidence intervals. Experimental results demonstrated that this method achieved the lowest calibration error (ECE reduced from 4.06% to 0.31%), outperforming common techniques such as temperature scaling and Platt scaling.
Notably, research on calibrating confidence scores from deterministic models into well-calibrated probability predictions remains limited in deep learning-based wildfire risk studies. This may be attributed to the growing preference for probabilistic deep learning methods—such as Bayesian neural networks—that inherently estimate predictive distributions. Compared to deterministic models, probabilistic approaches typically exhibit superior calibration performance. They yield outputs that more accurately approximate true event probabilities while simultaneously quantifying both data-related (aleatoric) and model-related (epistemic) uncertainties. Related research on uncertainty estimation will be discussed in Section 7.1.3.
@@ 6.2.1 Recurrent Neural Networks
As discussed in previous sections, the occurrence and spread of wildfires are influenced by various factors, including meteorological conditions, fuel status, and historical wildfire data. These factors exhibit cumulative effects (Flannigan et al., 2013). Therefore, utilizing time series data enables models to capture the gradual changes in influencing factors, thus facilitating the prediction of future wildfire risks. The most classical methods in time series prediction tasks are RNNs, which include RNN, LSTM, and GRU.
RNN is a basic neural network structure with recurrent connections, designed to process sequential data. RNN integrates input from the current time step with output from the previous time step to model and predict sequence data. A simple RNN structure consists of an input layer, hidden layer, and output layer. Input data are time series data, i.e., feature vectors with dimensions $f\times d$. The hidden layer includes several hidden units, where each hidden unit outputs $\mathbb{h}_{t-m}$, the ’memory’ at any intermediate time $t-m(m<n)$, as a weighted sum of its previous ’memory’ and current time step’s input features:
@@
Classic RNN networks and their variants have been employed in wildfire risk prediction. For instance, Cheng and Wang (2008) designed a spatiotemporal Elman RNN-based prediction framework integrating historical wildfire and weather data to predict the annual average burned area in Canadian forests. Experiments demonstrated a prediction error of less than 0.5ha on areas of 500ha, validating the feasibility of using RNN for fire prediction. Similarly, a comparative study of ten ANNs based on meteorological data for predicting wildfire incidence in Australia found that Elman RNN exhibited the best performance, with average accuracy, sensitivity, and specificity exceeding 93% (Dutta et al., 2013). Furthermore, a network composed of a dynamic autoencoder and RNN was used to predict the next month’s burned area in five U.S. regions, where the dynamic autoencoder was employed to transform multidimensional time series data into RNN input features, and dual decoders were used for predicting fire spot likelihood over different time spans, i.e., one week and one month. Experiments showed that their method was slightly superior to comparative methods, generative network, and GRU (Chavalithumrong et al., 2021).
@@
While Jain et al. (2020) and others have highlighted the limited use of classical RNNs, our findings are consistent with these observations. The underutilization of RNNs may be attributed to their difficulties in handling long sequence data, where they are prone to gradient vanishing or exploding, making it challenging to capture long-term dependencies (Hochreiter, 1997; Gers et al., 2000). To address these issues, LSTM networks were developed. LSTMs incorporate three gating units: input gate, forget gate, and output gate, which control the flow and retention of information. These gates have trainable weights and decide whether to pass, forget, or output information based on the current input and the output from the previous time steps. This gating mechanism effectively resolves the issues of vanishing and exploding gradients, thereby enabling better handling of long sequences and capturing of long-term dependencies. Consequently, LSTMs are more prevalently applied in wildfire risk prediction than classical RNNs.
@@
Moreover, some studies use only historical wildfire occurrence data combined with LSTM to predict future regional wildfire risks. For example, Kadir et al. (2023) used historical MODIS wildfire data combined with LSTM to predict the spatiotemporal distribution of wildfires in Indonesia for the following year, with a success rate exceeding 90% and an error rate of only 6.94%. Notably, to reduce the amount of data, this study grouped several days’ worth of data into single days for training and testing. Furthermore, Hu et al. (2023a) combined LSTM with autoencoders, using historical wildfire data from 1992-2018 to predict the level of wildfire occurrence in high-risk areas such as California. The autoencoder was employed to generalize wildfire events from the overall dataset, enhancing the accuracy of anomaly event predictions through reconstruction errors.
Similar to the LSTM, the GRU addresses the challenges of gradient vanishing and explosion inherent in traditional RNNs through its gating mechanisms. GRU reduces the number of gating units found in LSTM, containing only a reset gate and an update gate which simplifies the model architecture and enhances computational efficiency (Chung et al., 2014). Despite its simpler configuration, GRU retains the capability to model long-term dependencies with performance comparable to that of LSTM (Jin et al., 2020a).
Studies such as Dzulhijjah et al. (2023) have compared the performance of LSTM, BiLSTM, CNN-LSTM, Stacked-LSTM, and GRU in predicting wildfire risk in the Indonesian region of Kalimantan, finding that GRU slightly outperformed the other models, though the advantage is not significant. In contrast, Gopu et al. (2023) observed better results with LSTM in a comparative study of wildfire prediction in the Montesinho Natural Park, Portugal. Additionally, Chavalithumrong et al. (2021) found that models integrating a dynamic autoencoder with RNN outperformed standard GRU models, suggesting that enhanced RNN architectures can offer superior performance.
@@ 6.2.2 Transformers for Wildfire Risk Prediction
where $d_{k}$ represents the dimensionality of the key. The architecture of a Transformer comprises encoders, decoders, positional encoding, and multiple attention mechanisms.
@@
Transformers been adapted for several predictive tasks in wildfire risk prediction. For example, Miao et al. (2023) developed a Transformer model with a window-based attention mechanism to predict forest fire risks in Chongli District, Beijing, China, using time-series data of meteorological, topographical, vegetation, and anthropogenic factors. The window mechanism confines attention to a fixed-size window centered on each element, reducing computational complexity and memory usage. Similarly, Cao et al. (2024) employed a Transformer to predict next-day wildfire risks in Quanzhou County, Guangxi Province, China, based on three days of meteorological, topographical, and human activity data. The model’s performance, analyzed using IGR, outperformed LSTM, RNN, and SVM in terms of generalization, noise resistance, and overfitting mitigation.
@@ 6.3.1 Convolutional Neural Networks
CNNs are foundational network structures in deep learning, designed to extract local spatial features through convolution operations, making them well-suited for capturing the spatial information necessary for wildfire risk prediction, as shown in Fig 4. For example, Zhang et al. (2019) employed a modified AlexNet (Krizhevsky et al., 2012) using annual fire point maps and springtime mean influencing factors for Yunnan Province, China, to classify image patches as wildfire-prone areas. The model outputs a probability map of wildfire risk through softmax mapping in the final CNN layer. When comparing the CNN model’s performance with RF, SVM, MLP, and kernel logistic regression models, the authors found that CNN achieved the highest overall accuracy. They also analyzed the contributions of various wildfire drivers using the information gain ratio, identifying temperature, wind speed, surface roughness, precipitation, and elevation as the most influential factors.
@@
Building upon this, the authors extended the modified AlexNet to generate a global quarterly wildfire susceptibility map (Zhang et al., 2021a). They compared its performance with MLP-2D, CNN-1D, and MLP-1D models, finding that CNN-2D and MLP-1D models outperformed CNN-1D and MLP-2D. Additionally, they applied an improved permutation importance (PI) method and partial dependence plots (PDPs) to investigate the interpretability of the CNN-2D model. The PI analysis highlighted that maximum monthly temperature, soil temperature, NDVI, and soil moisture were significant factors affecting the model. Furthermore, PDP analysis revealed a negative relationship between cumulative precipitation and wildfire probability, while maximum monthly temperature and soil temperature positively influenced wildfire occurrence.
@@
Moreover, using a similar CNN approach, they predicted the spatiotemporal variation of wildfire risk under different climate scenarios, discovering that global warming would lead to an increase in burned areas and a northward shift of wildfire-prone regions (Zhang et al., 2024a). In another study, Kanwal et al. (2023) explored CNN-1D and CNN-2D classification models along with other machine learning algorithms for evaluating seasonal wildfire risks in Pakistan, finding that CNN-2D outperformed CNN-1D and other machine learning algorithms.
@@
In addition, some studies have employed deeper or structurally optimized CNN networks. For instance, Oak et al. (2024) utilized VGG16 with ReLU, BatchNorm, Dropout, and a sigmoid layer to predict wildfire susceptibility in Quebec, Canada. Comparative experiments demonstrated that this model outperformed other architectures like Xception, VGG16, ResNet50, and Inception V3. Similarly, Nur et al. (2022) applied a meta-heuristic optimization algorithm to fine-tune the hyperparameters of a CNN model for wildfire susceptibility mapping in the Plumas National Forest, USA. Their IGR analysis revealed that land use, drought index, and maximum temperature had the greatest impact on wildfire susceptibility, and the model’s performance improved significantly after optimization.
@@
Beyond classification-based methods, some studies have implemented end-to-end models for higher-resolution predictions. For example, Hodges and Lattimer (2019) proposed a deep convolutional inverse graphic network comprising convolution and deconvolution layers to simulate wildfire spread over homogeneous and heterogeneous landscapes for 6-hour and 24-hour intervals, demonstrating the high efficiency of CNNs in wildfire spread prediction. Similarly, Santopaolo et al. (2021) employed an end-to-end CNN-based image segmentation model to predict wildfire risk in Sicily, Italy, and California, USA, using the average values of driving factors over the preceding eight days.
@@
Furthermore, Allaire et al. (2021) developed a hybrid architecture combining convolutional and fully connected layers to emulate the output of a numerical fire spread simulator based on the Rothermel model of fire propagation. Similarly, Bergado et al. (2021) designed an AllConvNet, a fully convolutional network, to predict wildfire burn probability for the next 7 days in Victoria, Australia. This model achieved higher prediction accuracy compared to logistic regression, SegNet, and MLP models. Feature extraction and normalization of logistic regression coefficients revealed that total rainfall, lightning density, and surface temperature were among the most influential variables.
@@
Additionally, Li et al. (2023c) introduced spatial and channel attention mechanisms to enhance an end-to-end CNN model’s focus on critical features while suppressing irrelevant ones for predicting the next moment’s wildfire spread. This model also incorporated the burning map from the previous time step, outperforming other CNN models and ConvLSTM. Similarly, Shadrin et al. (2024) used the MA-Net segmentation model, which includes position attention and multi-scale attention blocks, to predict wildfire extent for the next 1-5 days. Unlike previous studies, they included short-term weather forecasts along with various wildfire drivers. Experimental results showed that their model accurately predicted wildfire spread within 3 days, with the wind components, land cover, NDVI, EVI, and LAI being the most influential factors.